OpenStudy (anonymous):

A hospital sends an invoice to a patient. The patient schedules a payment plan in which she makes an initial payment of $1,451 the first month and then pays one-third of the previous month’s payment each of the following months until the invoice is paid in full. Approximately how much of the invoice will be paid after the fourth month?

OpenStudy (texaschic101):

1451 -- first months payment 1/3(1451) = 1451/3 = 483.67 -- second month payment 1/3(483.67) = 483.67/3 = 161.22 -- 3rd month payment 1/3(161.22) = 161.22/3 = 53.74 -- 4th month payment 1451 + 483.67 + 161.22 + 53.74 = 2149.63 will be paid after the 4th month I am not sure if this is the answer you are looking for.